OpenStudy (anonymous):

How many liters of 10 M barium chloride solution are required to react with 50 mL of 18 M sulfuric acid solution? BaCl2 + H2SO4 → BaSO4 + HCl

OpenStudy (aaronq):

okay, is the equation balanced?

OpenStudy (anonymous):

um yes i think?

OpenStudy (aaronq):

hmm not quite, look at the number of Cl's and H's

OpenStudy (anonymous):

oh so u need a 2 in front of hcl?

OpenStudy (anonymous):

can u hang on I'm gunna reload the page it froze

OpenStudy (aaronq):

yeah no problem.

OpenStudy (aaronq):

thats right. okay, so in all these stoichiometry problems, the first step is to convert what's given to moles, ALWAYS (there may be some exceptions). So convert 50 mL of 18 M H2SO4 to moles. Hint: use the molarity formula

OpenStudy (anonymous):

I'm confused?

OpenStudy (aaronq):

you need to find the amount of moles of sulfuric acid. \(Molarity=\dfrac{moles~of~solute}{Liters ~of~{solution}}\)

OpenStudy (anonymous):

then?

OpenStudy (aaronq):

then you use a ratio, with the moles and the stoichiometric coefficients: \(\dfrac{moles~of~sulfuric~acid}{1}=\dfrac{moles~of~barium~chloride}{1}\) once you find the moles of barium chloride, you use the molarity formula again, but with the values for barium chloride. \(Molarity~of~BaCl_2=\dfrac{moles~of~barium~chloride}{Liters~of~solution}\) solve for liters of solution